OpenStudy (austinl):

Solve the initial value problem and determine where the solution attains it's minimum value. \(y\prime=2y^2+xy^2,~y(0)=1\)

OpenStudy (anonymous):

isn't this seperable?

OpenStudy (anonymous):

\[ \frac{dy}{dx} = 2y^{2} + xy^{2} \text{, }y(0)=1 \] \[\frac{dy}{dx} = y^{2}(2 + x)\Rightarrow \frac{dy}{y^{2}} = (2 + x)\, dx\]

OpenStudy (psymon):

Yeah, separable. Looks like no issue from there.

OpenStudy (austinl):

Why do I always forget this?? Thanks for the help!

OpenStudy (psymon):

Because we still gotta get used to doing it xD

OpenStudy (anonymous):

check with you subconscious...

OpenStudy (psymon):

I try to, but then it responds in russian.

OpenStudy (anonymous):

do you speak russian?

OpenStudy (psymon):

Roshia-go hanasemasen. (no IME on school computers -_-)

OpenStudy (anonymous):

i'm lost...

OpenStudy (psymon):

No, I cant. I take Japanese, not Russian, haha

OpenStudy (austinl):

привет

OpenStudy (anonymous):

darn subconscious... always making everything more difficult!

OpenStudy (psymon):

Yo quiero Taco Bell?

OpenStudy (anonymous):

talking to yourself again, i see...

OpenStudy (austinl):

Ok, back on subject :P

OpenStudy (austinl):

I have arrived at, \(\large{\frac{1}{y}=\frac{x^2}{2}+2x-1}\) How would I proceed from here in regards to the problem?

OpenStudy (loser66):

your solution is not right

OpenStudy (austinl):

\(\large{-\frac{1}{y}}\)

OpenStudy (loser66):

it should be \[-\frac{1}{y}=\frac{x^2}{2}+2x +C= \frac{x^2+4x+2C}{2}\\\frac{1}{y}=\frac{x^2+4x+2C}{-2} \]now, flip it up you have y = \(\dfrac{-2}{x^2+4x +2C}\) replace y(0) =1 to find out C you have C = -1 replace back to y , at that time, take limit of y you can see your solution

OpenStudy (austinl):

\(\large{y=-\dfrac{2}{x^2+4x-2}}\)

OpenStudy (loser66):

the denominator is an upward parabola, right? when it is bigger, y is smaller, right? so, it is bigger when x \(\rightarrow -\infty ~~ or +\infty, right?\)

OpenStudy (loser66):

and if x \(\pm \infty\) y \(\rightarrow 0\)

OpenStudy (austinl):

That is where the solution attains its minimum value? \(x\rightarrow \pm\infty\)?

OpenStudy (loser66):

oh, I may makes mistake, heehe, sorry, I forgot the minus sign in the front, I am so sorry, , so it turns your parabola is downward and its vertex is its maximum value . At that value, your y is minimum. You have to find out that vertex to plug it into your y.

OpenStudy (loser66):

got me?

OpenStudy (austinl):

Looking at these graphs, it approaches zero as x approaches infinity....

OpenStudy (loser66):

but you have local minimum at x =-2 still.

OpenStudy (austinl):

So is that the minimum value that it attains? \(x = -2\)

OpenStudy (loser66):

I think so.

OpenStudy (austinl):

Thanks very much @Loser66

OpenStudy (austinl):

And thank you too @pgpilot326 :)

OpenStudy (anonymous):

wait... if x->infinity, the whole thing goes to -infinity, right? that says no minimum

OpenStudy (anonymous):

likewise if x goes to -infinity...

OpenStudy (austinl):

\(\lim_{x \to \infty}~\dfrac{-2}{x^2+4x-1}=0\)

OpenStudy (anonymous):

sorry... brain fart

OpenStudy (austinl):

I dunno, I will just have to ask my professor.

OpenStudy (anonymous):

yeah but the bottom had 2 real solutions... \[2 \pm \sqrt{6}\]

OpenStudy (anonymous):

so the min will occur there and will be \[-\infty\]

OpenStudy (anonymous):

oops, \[-2 \pm \sqrt{6}\]

OpenStudy (anonymous):